# Making Sense of Gravitational Thermodynamics

### v1

@@ 4.1 Statistical physics of idealised self-gravitating systems
from which we can define $T=1/\beta$ (temperature) and $\mu=\alpha/\beta$ (Gibbs chemical potential) which are global thermodynamic quantities of the whole system. Notice that the partial derivatives must be evaluated at the extremum, which highlights the role of equilibrium states. Recall that one of the hallmarks of the problematic nature of lack of extensivity was the inability to derive equations like (3). We take the definability of these quantities and relations as another proof of the applicability of equilibrium statistical mechanics over and above non-equilibrium statistical mechanics. Furthermore, this also supports the claim that $N\rightarrow\infty$ should be taken as the salient limit to recover macroscopic quantities within SGS instead of the $N,V\rightarrow\infty$ (constant $N/V$) thermodynamic limit. The latter might be the correct limit to adopt for short-range interacting systems if we want to recover macroscopic quantities, but there is no reason to believe it is true for every kind of system and thus we should instead be open to employing different limits in other contexts if we want to recover those kinds of coarser-grained features. This can help to address Callender’s worry about the arbitrariness of adopting this limit. [^32] We return to the justification of this limit in §5. In that section we also explain why the $N\rightarrow\infty$ limit deserves the name of thermodynamic limit.
@@
    The arguments provided rely on idealised models. In particular, real clusters are evidently not in boxes, but the box idealisation has been adopted from the outset without clear justification. Similarly, we have so far considered only collisionless systems. These features are drastic departures from realistic physical examples of SGS like globular clusters. Hence we should ask whether they are justified, and what can we learn about actual systems from those descriptions.
@@
    This section supported the application of equilibrium statistical mechanics to SGS, but we were also able to recover certain quantities and relations that are often characterised as ‘thermodynamic’, and we employed what can be regarded as a ‘thermodynamic’ type of limit. We ask whether thermodynamics still has a place in the physics of SGS and which place can it have, especially since the arguments of this section countered one of Robertson’s main reasons against the applicability of thermodynamics (lack of stable equilibrium), and questioned the claim that the physics of SGS does not abstract away from microdetails.
@@ 4.2 Removing idealisations
We start by highlighting that the box idealization has clear effects on SGS models under certain conditions. If we model SGS as systems confined in a box that allows the stars to bounce against the walls without losing energy, then we can always make the system stable and in equilibrium by restricting the radius of the box enough (Heggie and Hut, 2003, p. 177). If the radius is small enough, kinetic energy dominates and the effect of gravity is negligible, thus the system behaves like an ideal gas and is not in virial equilibrium. As Heggie and Hut (2003) remark, systems in these models are stable, as proved by Lynden-Bell and Wood (1968). Thus, if the box is small enough the behaviour of the system is heavily influenced by it and equilibrium statistical mechanics ‘trivially’ applies. What about bigger boxes? [^34] For a more quantitative description of the scenario see Heggie and Hut (2003, §8, §17) and Padmanabhan (1990, §4.4). If the box is large enough then the stability depends less trivially on the energy of the system, as per Figures 1 and 2 above. If the energy is high enough, then as before the kinetic energy dominates over the gravitational effects and we have metastable quasi-equilibrium states as argued in the previous subsection. However, the idealisation is still influencing the model’s behaviour and thus it is reasonable to inquire about its justifiedness. Finally, in the isothermal models of §4.1 there was no mention of collisions and interactions between stars, and collisionless evolution naturally renders stability easier to secure. It is reasonable to ask whether taking into account interactions makes stability unattainable.
@@
In a nutshell, the King model is a spherically symmetric modified isothermal model. In this model, stars are not mechanically confined by any boundaries and are allowed to escape (although the system is still generally bound gravitationally). However, the system is still parametrised by a radius to avoid the problem of infinite mass and the divergence of the density of states: if stars that evaporate are still considered in the system, then the system’s extension will tend to infinity and the density of states will diverge making the statistical mechanical description not well-defined. To make the system statistically tractable, the model introduces a cutoff radius which is determined by the escape energy of the high-energy stars that evaporate (Meylan and Heggie (1997), Chavanis (2006)). The distribution function of the King model is:
@@
What happens to these systems now that the box is removed? We focus on globular clusters in the King model and present their evolution taking also into account the effects of collisions and interactions. These models behave like real globular clusters, and these astrophysical objects are believed to naturally tend towards core collapse in normal conditions once encounters take place. But let’s first distinguish two time scales, the crossing time $t_{cross}=R/v$ relating size of the system and velocity of stars and the relaxation time $t_{relax}\sim\frac{0.1N}{lnN}t_{cross}$, where $N$ is the number of stars in the system. The crossing time is the time needed for a typical star to cross the galaxy once and is “the shortest time scale on which the system as a whole can react to global changes in its potential”, the relaxation time is “the time scale in which the cumulative effect of two-body encounters can alter the individual stellar orbits significantly” (Elson et al., 1987, p. 566). The relaxation time is significant for globular clusters and on this time scale encounters drive the evolution of stellar systems by different mechanisms (Meylan and Heggie (1997), Binney and Tremaine (2011, §7)).
@@
Hence the system naturally slowly evolves towards a core-halo structure with a progressively shrinking core, and overall there is a decrease in the radius of the cluster and an increase in the concentration parameter $c$. For $c<2.1$ the cluster follows the King sequence closely, whereas for larger values of $c$ the King models become unstable. [^37] As Elson et al. remark: “This is in reasonable agreement with observations of most globular clusters, as well as with numerical calculations of globular cluster evolution in the Fokker-Planck approximation, which show that an evolving cluster follows the King sequence closely as long as $c<2.1$(Elson et al., 1987, pp. 568-9). This instability, powered by the negative heat capacity, sets the final stage of the core collapse in the form of gravothermal (or isothermal) collapse (Chavanis, 2006, §4.9-4.10). [^38] See also Chavanis et al. (2015, p. 5) for a description of self-gravitating systems within the King model (i.e. without the box). It should be noted that the same phenomenon happens in the box-confined idealised case when the box is big enough that kinetic energy does not dominate and thus it becomes entropically favourable to form a core-halo structure. Following this runaway process, the system collapses up to a point in which binary stars’ interactions emit enough energy to make the system re-expand and this triggers gravothermal oscillations. [^39] See Heggie and Hut (2003, Part VIII) on post-collapse evolution. Two classic works on the topic are Cohn and Hut (1984) and Hut (1983).
@@
Figure 3: Series of equilibria $(E,\beta)$ for different kinds of models. (Katz, 1980)
@@
Before moving on, a note on the role of negative heat capacity within the statistical physics of SGS is in order. As pointed out in §3.2, Callender (2011) asks whether negative specific heat would ultimately be unavoidably problematic for equilibrium and whether negative specific heat can genuinely hold in equilibrium systems. The description of SGS presented in this section dispels these worries. Core collapse and negative heat capacity are genuinely displayed by SGS but they only happen in certain regimes. Hence they can still be described as evolving through a series of effective equilibria in the right regime even if negative heat capacity can be instantiated at some point. This unconventional aspect of SGS further proves how studying these systems allows us to learn interesting lessons about equilibrium statistical mechanics.
@@ 5 Minimal Thermodynamics
The previous sections have defended the applicability of equilibrium and stability within SGS and the scale at which they hold, and have also shown how coarse-grained quantities can be defined. This supported the applicability of the core scaffolding of equilibrium statistical mechanics to SGS. However, equilibrium statistical mechanics is obviously compatible with thermodynamics and indeed thermodynamics is usually taken to emerge from it. And, in the last section, we have been able to define within the equilibrium statistical mechanics of SGS quantities, relations and a limit that are usually regarded as ‘thermodynamic’. Here we return to the central question of the whole debate and ask whether thermodynamics does apply to SGS and, if yes, in what sense. We propose a minimal framework for thermodynamics and argue that it applies to the physics of SGS, vindicating a notion of gravitational thermodynamics.
@@
Briefly put, our thesis is that a minimal notion of thermodynamics applies to SGS because we can model them via macro-quantities and we can use these quantities to deliver higher-level non-statistical explanations. Minimal thermodynamics is not an additional theory but rather an additional level of description that can be embedded in the theoretical framework of equilibrium statistical mechanics. The quantities of the minimal thermodynamic description are coarse-grained quantities like energy, temperature and entropy that can be derived within equilibrium statistical mechanics but can enter higher-level explanations of the behaviour of SGS above the mere lower-level statistical mechanical behaviour expressed in terms of micro-variables and distribution functions. At the same time, we argue that minimal thermodynamics is not at the same level of description as phenomenological thermodynamics, which is even more abstract. Hence ‘minimal thermodynamics’ captures a distinctive level of description and layer of explanations. Here is the framework in a nutshell:
@@
    Autonomy: Minimal thermodynamics is partly autonomous from the statistical lower-level description, although it is based on quantities derived from it. Minimal thermodynamics concerns the evolution of the macroscopic quantities irrespectively from their statistical underpinning.
@@
The following table broadly illustrates the relationship between theories and levels of description in the domain of statistical physics and thermodynamics:
@@
Phenomenological Thermodynamics
Empirical descriptions in terms of macroscopic transformations on systems
$\delta Q$, $\delta W$, ...
Equilibr Stat Mech
Minimal Thermodynamics: Coarse-grained state functions and relations, etc.
$E$, $T$, $V$, $P$, $S$, ...
Equilibrium Stat Mech: Canonical and microcanonical ensemble, density of state, partition function, etc.
$g(E)$, $Z(\beta)$, ...
Non-Equilibrium Stat Mech
Evolution of one-particle distribution functions via Boltzmann equation, etc.
$f(\textbf{x},\textbf{p},t)$, ...
@@
A specific and clear example of the application of minimal thermodynamics is the descriptions and explanations of SGS based on the diagrams in Figures 1 and 2. Katz (1978, 1980, 2003) developed Poincaré’s linear series of equilibria theory to analyse the stability of SGS in terms of quantities such as $E$ and $\beta$ [^44] Following the approach by Antonov (1962) and Lynden-Bell and Wood (1968). Basically, turning points in the $(E,\beta)$ phase diagram denotes instabilities:
@@
Thus, looking at macro-level quantities provides a platform for developing an explanation of the stability of SGS and their evolution into a core-halo structure from a distinctive higher-level perspective. As stressed by Chavanis (2006, §4.4) and Katz (2003, §2.4, §3.3) the results we obtain with these methods are compatible with the results obtained via kinetic theory and numerical methods and thus are in principle derivable also from that more detailed lower-level perspective. Hence we have at least two distinct dimensions of descriptions and explanations working at the same time, one characteristically statistical mechanical and the other thermodynamic, in the sense of being coarse-grained with respect to micro-variables and e.g. distribution functions. [^45] More specifically, as mentioned in §4.1, the statistical-based dimension of explanation comprises both non-equilibrium and equilibrium statistical explanations. This shows how Katz’s approach adheres to points (I)-(II) of minimal thermodynamics.
Concerning (III), we can point out that these higher-level descriptions and explanations are partly autonomous from the statistical level: they are based on a diagram obtained by maximizing statistical entropy and so the $(E,\beta)$ description is derived from the statistical mechanical description, but Katz’s explanation of instability is purely expressed in terms of the behaviour of these macro-quantities independently from their underpinning. Relating this point to the literature on inter-level relations in science, this type of independence from the micro-level could be framed in terms of emergent behaviour, in a sense of emergence compatible with reduction, as these descriptions are in principle derivable. [^46] See for instance Butterfield (2011) and Franklin and Knox (2018). It could be classified as ‘coarse-grained emergence’:
A coarse-grained description of a system emerges synchronically upon a fine-grained description, iff the former has terms denoting properties or behavior that are novel and autonomous with respect to the latter, and these properties or behavior supervene upon the behavior of the components of the fine-grained description. (Palacios, 2022, p. 39)
@@
The approach just discussed is a specific example, but is representative of the general fact that in certain regimes SGS can be expressed in terms of macro-level quantities as presented in the previous section, for instance talking of virial equilibrium (as $02K+U=$ 0), negative heat capacity (as $C_{V}=(\partial E/\partial T)|_{V}<0$), temperature ($T^{-1}(E)=\beta(E)=\partial S(E)/\partial E$), and relations like $dE=TdS-P_{b}dV+\mu dN$. This relation derived within equilibrium statistical mechanics from the base micro-level description can be regarded as an alternative statement of the first law of thermodynamics, although this is not the general first law of thermodynamics in terms of work and heat introduced by phenomenological thermodynamics independently from any micro-description, i.e. $dU=d\bar{}\hskip 1.00006ptQ+d\bar{}\hskip 1.00006ptW$. This is thus a good example of a relation that is not fully thermodynamic in the sense of phenomenological thermodynamics but is also not based on micro-variables, and supports (I). Hence minimal thermodynamics captures a middle ground description between the two levels.
@@
All these things being considered, we highlight how the analysis of this section makes clear how the theory of phenomenological thermodynamics is not suitable for SGS. If we think of thermodynamics in those terms then it is right to conclude that thermodynamics does not apply to SGS as Robertson argues. Concepts like thermal contact, work, heat and the idea of gradually changing external parameters and performing operations on these systems make little sense. [^47] See Heggie and Hut (2003, p. 19) on the lack of handles we have on modelling SGS. Instead, the kind of thermodynamic descriptions we can develop in the physics of SGS match more with the type of thermodynamic relations we can formulate in equilibrium statistical mechanics. These relations are accommodated by minimal thermodynamics but are not part of phenomenological thermodynamics. For instance, as we mentioned, the relation $dE=TdS-P_{b}dV+\mu dN$ that can hold for SGS is the specific version of the first law of thermodynamics that is obtained in equilibrium statistical mechanics. That relation can in principle be used to derive the first law of thermodynamics $dU=d\bar{}\hskip 1.00006ptQ+d\bar{}\hskip 1.00006ptW$ as found in phenomenological thermodynamics, but they are not the same identical law, as one is empirically formulated and the other is derived within equilibrium statistical mechanics and they are based on different quantities. Furthermore, focusing on $T$, we stress that this is the temperature defined via statistical equilibrium and not temperature as introduced by thermodynamic equilibrium in phenomenological thermodynamics via the Zeroth law of thermodynamics.
@@
Wrapping up, there is space for thermodynamics in SGS physics even though the theory of phenomenological thermodynamics does not apply. SGS provided an ideal case study to develop this original take on thermodynamics. This minimal framework for thermodynamics is the most perspicuous and useful way to think about thermodynamics in SGS especially as it is grounded in physics practice. Also, describing minimal thermodynamics as a kind of thermodynamics vindicates the ‘thermodynamic’ label used by physicists in talking about thermodynamic methods in statistical SGS physics. [^48] For instance Chavanis (2006) distinguished between the thermodynamic approach used by Lynden-Bell and Wood (1968), Padmanabhan (1990) and himself and the statistical mechanical approach starting by the density of states or partition functions. We vindicate this classification but we reframe it in clearer terms. More broadly, adopting this framework clarifies how kinetic theory, equilibrium statistical mechanics, and a kind of thermodynamics embedded within the theoretical frame of equilibrium statistical mechanics are all employed at different levels in the physics of SGS. Every one of these three levels is equally justified and interpreting the physics of SGS leaving out one of these dimensions of description and explanation would miss something about our understanding of these systems.
6 Conclusion

### v2

@@ 4.1 Statistical physics of idealised self-gravitating systems
from which we can define $T=1/\beta$ (temperature) and $\mu=\alpha/\beta$ (Gibbs chemical potential) which are global thermodynamic quantities of the whole system. Notice that the partial derivatives must be evaluated at the extremum, which highlights the role of equilibrium states. We take the definability of these quantities and relations as another proof of the applicability of equilibrium statistical mechanics over and above non-equilibrium statistical mechanics. Furthermore, this also supports the claim that $N\rightarrow\infty$ should be taken as the salient limit to recover macroscopic quantities within SGS instead of the $N,V\rightarrow\infty$ (constant $N/V$) thermodynamic limit. The latter might be the correct limit to adopt for short-range interacting systems if we want to recover macroscopic quantities, but there is no reason to believe it is true for every kind of system and thus we should instead be open to employing different limits in other contexts if we want to recover those kinds of coarser-grained features. This can help to address Callender’s worry about the arbitrariness of adopting this limit. [^31] We return to the justification of this limit in §5. In that section we also explain why the $N\rightarrow\infty$ limit deserves the name of thermodynamic limit.
@@
    The arguments provided rely on idealised models. In particular, real clusters are evidently not in boxes, but the box idealisation has been adopted from the outset without clear justification. Similarly, we have so far considered only collisionless systems. These features are drastic departures from realistic physical examples of SGS like globular clusters. Hence we ask whether they are justified, and what can we learn about actual systems from those descriptions.
@@
    This section supported the application of equilibrium statistical mechanics to SGS. Hence it is reasonable to ask whether thermodynamics still has a place in the physics of SGS, especially since we countered Robertson’s main reasons against the applicability of thermodynamics.
@@ 4.2 Removing idealisations
We start by highlighting that the box idealization has clear effects on SGS models under certain conditions. If we model SGS as systems confined in a box that allows the stars to bounce against the walls without losing energy, then we can always make the system stable and in equilibrium by restricting the radius of the box enough (Heggie and Hut, 2003, p. 177). If the radius is small enough, kinetic energy dominates and the effect of gravity is negligible, thus the system behaves like an ideal gas and is not in virial equilibrium (and the heat capacity is positive). As Heggie and Hut (2003) remark, systems in these models are stable, as proved by Lynden-Bell and Wood (1968). Thus, if the box is small enough the behaviour of the system is heavily influenced by it and equilibrium statistical mechanics ‘trivially’ applies. What about bigger boxes? [^33] For a more quantitative description of the scenario see Heggie and Hut (2003, §8, §17) and Padmanabhan (1990, §4.4). If the box is large enough then the stability depends less trivially on the energy of the system, as per Figures 1 and 2 above. If the energy is high enough, then as before the kinetic energy dominates over the gravitational effects and we have metastable quasi-equilibrium states as argued in the previous subsection. However, the idealisation is still influencing the model’s behaviour and thus it is reasonable to inquire about its justifiedness. Finally, in the isothermal models of §4.1 there was no mention of collisions and interactions between stars, and collisionless evolution naturally renders stability easier to secure. It is reasonable to ask whether taking into account interactions makes stability unattainable.
@@
In a nutshell, the King model is a spherically symmetric modified isothermal model. In this model, stars are not mechanically confined by any boundaries and are allowed to escape (although the system is still generally bound gravitationally). However, the system is still parametrised by a radius to avoid the problem of infinite mass and the divergence of the density of states: if stars that evaporate are still considered in the system, then the system’s extension will tend to infinity and the density of states will diverge making the statistical mechanical description not well-defined. To make the system statistically tractable, the model introduces a cutoff radius which is determined by the escape energy of the high-energy stars that evaporate (Meylan and Heggie (1997), Chavanis (2006b)). The distribution function of the King model is:
@@
What happens to these systems now that the box is removed? We focus on globular clusters in the King model and present their evolution taking also into account the effects of collisions and interactions. [^34] In the context of globular clusters, by collision we mean close gravitational encounters. These models behave like real globular clusters, and these astrophysical objects are believed to naturally tend towards core collapse in normal conditions once encounters take place. But let’s first distinguish two time scales, the crossing time $t_{cross}=R/v$ relating size of the system and velocity of stars and the relaxation time $t_{relax}\sim\frac{0.1N}{lnN}t_{cross}$, where $N$ is the number of stars in the system. The crossing time is the time needed for a typical star to cross the galaxy once and is “the shortest time scale on which the system as a whole can react to global changes in its potential”, the relaxation time is “the time scale in which the cumulative effect of two-body encounters can alter the individual stellar orbits significantly” (Elson et al., 1987, p. 566). The relaxation time is significant for globular clusters and on this time scale encounters drive the evolution of stellar systems by different mechanisms (Meylan and Heggie (1997), Binney and Tremaine (2011, §7)).
@@
Hence the system naturally slowly evolves towards a core-halo structure with a progressively shrinking core, and overall there is a decrease in the radius of the cluster and an increase in the concentration parameter $c$. For $c<2.1$ the cluster follows the King sequence closely, whereas for larger values of $c$ the King models become unstable. [^37] As Elson et al. remark: “This is in reasonable agreement with observations of most globular clusters, as well as with numerical calculations of globular cluster evolution in the Fokker-Planck approximation, which show that an evolving cluster follows the King sequence closely as long as $c<2.1$(Elson et al., 1987, pp. 568-9). This instability, powered by the negative heat capacity, sets the final stage of the core collapse in the form of gravothermal (or isothermal) collapse (Chavanis, 2006b, §4.9-4.10). [^38] See also Chavanis et al. (2015, p. 5) for a description of self-gravitating systems within the King model (i.e. without the box). It should be noted that the same phenomenon happens in the box-confined idealised case when the box is big enough that kinetic energy does not dominate and thus it becomes entropically favourable to form a core-halo structure. Following this runaway process, the system collapses up to a point in which binary stars’ interactions emit enough energy to make the system re-expand and this triggers gravothermal oscillations. [^39] See Heggie and Hut (2003, Part VIII) on post-collapse evolution. Two classic works on the topic are Cohn and Hut (1984) and Hut (1983).
@@
Figure 3: Series of equilibria $(E,\beta)$ for different models. (Katz, 1980, Figure 1)
@@
Before moving on, a note on the role of negative heat capacity within the statistical physics of SGS is in order. As pointed out in §3.2, Callender (2011) asks whether negative specific heat would ultimately be unavoidably problematic for equilibrium and whether negative specific heat can genuinely hold in equilibrium systems. This section dispelled these worries. Core collapse and negative heat capacity are genuinely displayed by SGS, but only in certain regimes, and only under certain conditions negative heat capacity makes the system completely unstable. Hence SGS can still be described as evolving through a series of effective equilibria in the right regime even if negative heat capacity can be instantiated at certain points.
@@ 5 Minimal Thermodynamics
The previous sections have supported the applicability of the core scaffolding of equilibrium statistical mechanics to SGS. However, equilibrium statistical mechanics is compatible with thermodynamics and indeed thermodynamics is usually taken to emerge from it. And, in the last section, we have been able to define within the equilibrium statistical mechanics of SGS quantities, relations and a limit that are usually regarded as ‘thermodynamic’. Here we return to the central question of the whole debate and ask whether thermodynamics does apply to SGS and, if yes, in what sense. Based on the results of the previous sections, we argue that a ‘minimal’ thermodynamic level of description is available in the case of SGS, even if the phenomenological thermodynamics level of description in terms of e.g. heat, work, and Clausius’ second law of thermodynamics expressed via heat transfer has no practical use here.
In a nutshell, we argue that we can devise macroscopic descriptions and models of SGS as well as explanations of their behaviour via macroscopic quantities like state functions in terms of (e.g.) energy and temperature over and above the purely statistical descriptions in terms of (e.g.) partition functions. We call them minimal thermodynamics descriptions and explanations: ‘thermodynamics’ because they involve macroscopic descriptions of these systems above the level of description of the partition function and density of states, ‘minimal’ because this is still not the level of description of phenomenological thermodynamics.
This middle-way level of description abstracts away from the statistical descriptions of SGS by introducing new macroscopic quantities allowing for simpler novel explanations of the behaviour of SGS, largely autonomous from the lower-level more complex statistical level. This provides a new perspicuous conceptual framing of the physics of SGS and a finer-grained account of all the levels at which it operates, and vindicates a notion of gravitational thermodynamics. There are four levels of descriptions within statistical and thermal physics, corresponding to phenomenological thermodynamics, minimal thermodynamics, equilibrium statistical mechanics, and non-equilibrium statistical mechanics. Based on §4 we say that only the last three apply to SGS. These are distinct levels and we can provide different types of explanations of the behaviour of SGS at each level. Merging together two or more of these levels would render our understanding of the physics of SGS less fine-grained.
5.1 Introducing minimal thermodynamics
We introduce the notion of minimal thermodynamics in general terms. The next subsections clarify the framework by applying it to SGS. Minimal thermodynamics refers to a descriptive framework situated between statistical mechanics, which uses ensembles, and phenomenological thermodynamics, which focuses e.g. on heat exchange. We frame minimal thermodynamics not as an additional theory but rather as an additional level of description that can be embedded in the theoretical framework of equilibrium statistical mechanics. [^44] Although we are not committed to a strict distinction between theories and levels of description, an intuitive way to see how they differ is by noticing how we can have different levels of description within the framework of one theory. For example, the same system in the same theory can be described at different levels of abstraction (Franklin and Knox, 2018). We include it there and not within phenomenological thermodynamics because we understand the latter as a ‘control’ theory focused on operations we can do on and between systems involving concepts like heat exchange, work, or Carnot cycles. [^45] However, this is mostly a semantic point and our presentation does not hinge on that: in principle, one would be free to follow Callender (2011) and relax our concept of phenomenological thermodynamics and argue that minimal thermodynamics is another lower level of description within the theory of phenomenological thermodynamics in addition to the traditional level of phenomenological thermodynamics just described. Hence we stress that our approach is consistent with Callender’s proposal of relaxing our notion of thermodynamics.
@@ 5.1 Introducing minimal thermodynamics
The quantities of the minimal thermodynamic description are coarse-grained quantities like energy, temperature and entropy that can be derived within equilibrium statistical mechanics but can enter higher-level explanations of the behaviour of SGS above the mere lower-level statistical mechanical behaviour expressed in terms of micro-variables and distribution functions. At the same time, minimal thermodynamics is not at the same level of description as phenomenological thermodynamic descriptions, which are even more abstract. Hence ‘minimal thermodynamics’ captures a distinctive level of description and layer of explanations. We stress that this notion is meant to be general and not restricted to the case of SGS. Here is the general framework (see also Table 1):
@@
    Autonomy and novelty: Minimal thermodynamics is partly autonomous from the statistical lower-level description, although it is based on quantities derived from it. Minimal thermodynamics concerns the evolution of the macroscopic quantities irrespective of their statistical underpinning. Minimal thermodynamic explanations can be novel insofar as the explanations we devise via the macroscopic variables above offer additional explanations of a given phenomenon which were not available in the same form at a lower level of description.
@@
The features of this framework align with the widespread trend in the philosophy of science that emphasizes the crucial role of abstractions in scientific explanations (Batterman (2002, 2009, 2021), Strevens (2008), Weslake (2010), Morrison (2015)). The general idea is that more abstract and less precise explanations can be nonetheless more powerful than more detailed explanations in certain respects, for instance in virtue of being more practically useful. The framework of minimal thermodynamics is especially consistent with the account of abstraction developed by Knox (2016, 2023) and Franklin and Knox (2018). They defend the usefulness of abstractions and in particular abstractions obtained by change of variables that allow us to screen off certain details. Such abstractions change the level of description at which a certain system is described, and enable a simpler representation of the system, facilitating novel and more effective explanations of the same phenomena. Their examples include changes of variables in the descriptions of harmonic oscillators, which lead to simpler equations, or—more interestingly—changes of variables that make the descriptions of phonons more perspicuous. In the case of SGS, we argue that introducing diagrams like those in Figures 1–3, where SGS are modelled solely in terms of energy and temperature, allows us to analyse certain aspects of their behaviour using the simple series of equilibria method. This approach provides simpler and useful models that are not available at lower, less abstract levels of description. This is the sense in which focusing levels of description is particularly interesting in the present context.
Table 1: Theories and levels of description.
@@
Phenomenological Thermodynamics Empirical descriptions in terms of macroscopic transformations on systems $\delta Q$, $\delta W$, ...
Equilibr Stat Mech Minimal Thermodynamics: Coarse-grained state functions and relations, etc. $E$, $T$, $V$, $P$, $S$, ...
Equilibrium Stat Mech: Canonical and microcanonical ensemble, density of state, partition function, etc. $g(E)$, $Z(\beta)$, ...
Non-Equilibrium Stat Mech Evolution of one-particle distribution functions via Boltzmann equation, etc. $f(\textbf{x},\textbf{p},t)$, ...
5.2 Minimal thermodynamics of self-gravitating systems
@@
A specific and clear example of the application of minimal thermodynamics is the descriptions and explanations of SGS based on the diagrams in Figures 1, 2 and 3. Katz (1978, 1980, 2003) developed Poincaré’s linear series of equilibria theory to analyse the stability of SGS in terms of quantities such as $E$ and $\beta$ [^46] Following the approach by Antonov (1962) and Lynden-Bell and Wood (1968). Basically, turning points in the $(E,\beta)$ phase diagram denotes instabilities:
@@
Thus, looking at macro-level quantities provides a platform for developing an explanation of the stability of SGS and their evolution into a core-halo structure from a distinctive higher-level perspective. As stressed by Chavanis (2006b, §4.4) and Katz (2003, §2.4, §3.3) the results we obtain with these methods are compatible with the results obtained via kinetic theory and numerical methods and thus are in principle derivable also from that more detailed lower-level perspective. Hence we have at least two distinct dimensions of descriptions and explanations working at the same time, one characteristically statistical mechanical and the other thermodynamic, in the sense of being coarse-grained with respect to micro-variables and e.g. distribution functions. [^47] More specifically, as mentioned in §4.1, the statistical-based dimension of explanation comprises both non-equilibrium and equilibrium statistical explanations. This shows how Katz’s approach adheres to points (I)-(II) of minimal thermodynamics.
Concerning (III), we point out that these higher-level descriptions and explanations are partly autonomous from the statistical level: they are based on a diagram obtained by maximizing statistical entropy and so the $(E,\beta)$ description is derived from the statistical mechanical description, but Katz’s explanation of instability is purely expressed in terms of the behaviour of these macro-quantities independently from their underpinning. Relatedly, these explanations can be novel in the sense of §5.1 for the reasons given in the previous paragraph: introducing the minimal thermodynamic description in terms of state variables allows for building new models of SGS and a novel dimension of explanations for SGS behaviour on top of the kinetic theory level and the ensemble level. Relating this point to the literature on inter-level relations in science, this type of independence from the micro-level and novelty could be framed as emergent behaviour, in a sense of emergence compatible with reduction, as these descriptions are in principle derivable. [^48] See for instance Butterfield (2011) and Franklin and Knox (2018). More precisely, it could be classified as ‘coarse-grained emergence’, following Palacios (2022, p. 39).
@@
The approach just discussed is a specific example, but is representative of the general fact that in certain regimes SGS can be expressed in terms of macro-level quantities as presented in the previous section. For instance, virial equilibrium (as $02K+U=$ 0), negative heat capacity (as $C_{V}=(\partial E/\partial T)|_{V}<0$), and temperature ($T^{-1}(E)=\beta(E)=\partial S(E)/\partial E$) are central to Lynden-Bell and Wood (1968) and the related literature. Also, as presented in §4.1, ‘thermodynamic’ functions in the sense of minimal thermodynamics are used by e.g. Tatekawa et al. (2005) and Casetti and Nardini (2010) within a variety of SGS models. Furthermore, relations like $dE=TdS-P_{b}dV+\mu dN$ are defined (Katz, 2003, §3.1). [^49] Although it has to be stressed that this relation is derived in the idealised situation described in §4.1. This relation derived within equilibrium statistical mechanics from the base micro-level description can be regarded as an alternative statement of the first law of thermodynamics, although this is not the general first law of thermodynamics in terms of work and heat introduced by phenomenological thermodynamics independently from any micro-description, i.e. $dU=d\bar{}\hskip 1.00006ptQ+d\bar{}\hskip 1.00006ptW$. This is thus a good example of a relation that is not fully thermodynamic in the sense of phenomenological thermodynamic descriptions but is also not based on micro-variables, and supports (I). Hence minimal thermodynamics captures a middle ground level of description between the two levels.
@@
All these things being considered, this section has clarified how the theory of phenomenological thermodynamics is unsuitable for SGS. If we think of thermodynamics purely in terms of phenomenological thermodynamics as we introduced it then it is reasonable to conclude that thermodynamics does not apply to SGS as Robertson concludes, by definition. Concepts such as thermal contact, work, heat, and the gradual variation of external parameters, along with operations performed on these systems, hold little relevance — or are entirely absent —from the actual physics literature on SGS. [^50] See Heggie and Hut (2003, p. 19) on the lack of handles we have on modelling SGS. On this, we agree with Robertson. Instead, the kind of ‘thermodynamic’ descriptions we can develop in the physics of SGS match more with the type of thermodynamic relations we can formulate in equilibrium statistical mechanics. These relations fall within the domain of minimal thermodynamic descriptions rather than phenomenological thermodynamic descriptions. For instance, as we mentioned, the relation $dE=TdS-P_{b}dV+\mu dN$ that holds for certain SGS models is the specific version of the first law of thermodynamics that is obtained in equilibrium statistical mechanics. That relation can in principle be used to derive the first law of thermodynamics $dU=d\bar{}\hskip 1.00006ptQ+d\bar{}\hskip 1.00006ptW$ as found in phenomenological thermodynamics, but they are not the same identical law, as one is empirically formulated and the other is derived within equilibrium statistical mechanics and they are based on different quantities. In this sense, they belong to two different levels of descriptions and abstraction, as articulated above. [^51] On this we agree with Franklin and Knox (2018) who hold that ‘higher-level’ descriptions obtained from abstractions and change of variables imposed on ‘lower-level’ descriptions can be regarded as novel and weakly emergent even if reduction is available. Whereas the level of description at which quantities like energy and temperature is available and employed within the physics of SGS, descriptions of SGS via heat exchange or work are absent from the physics literature on SGS (assuming, of course, that it is even meaningful to discuss concepts like heat exchange within isolated astrophysical systems). Furthermore, focusing on $T$, we stress that this is the temperature defined via statistical equilibrium and not temperature as introduced by thermodynamic equilibrium in phenomenological thermodynamics via the Zeroth law of thermodynamics, bolstering the claim that these are two descriptive levels.
@@
Wrapping up, there is still space for thermodynamics in SGS physics. The minimal framework for thermodynamics presented here is the most perspicuous way to think about thermodynamics in SGS especially as it is grounded in physics practice. Also, describing minimal thermodynamics as a kind of thermodynamics vindicates the ‘thermodynamic’ label used by physicists in talking about thermodynamic methods in statistical SGS physics. [^52] For instance Chavanis (2006b) distinguished between the thermodynamic approach used by Lynden-Bell and Wood (1968), Padmanabhan (1990) and himself and the statistical mechanical approach starting by the density of states or partition functions. More broadly, adopting this framework clarifies how kinetic theory, equilibrium statistical mechanics, and a kind of thermodynamics embedded within the theoretical frame of equilibrium statistical mechanics are all employed at different levels in the physics of SGS. Every one of these three levels is equally justified and interpreting the physics of SGS leaving out one of these dimensions of description and explanation would miss something about our understanding of these systems.
6 Conclusion and Future Directions